# Scattering of CO from vacant-MoSe2 with O adsorbates: is CO2 formed?

### v1

@@ 4 Conclusions
We have studied the interaction of CO molecules coming from vacuum and a 2O-MoSe2-vSe surface, in which one O atom is adsorbed in a Se vacancy and a second O atom is adsorbed on top of a Se atom of the phase of the MoSe2 monolayer. The latter is the equilibrium structure upon (dissociative) adsorption of a O2 molecule on a MoSe2 monolayer with one Se vacancy 51. Interestingly, DFT calculations show that the CO recombination with the O atom adsorbed on top of the Se atom with either subsequent adsorption or desorption of the formed CO2 molecules are highly exothermic reactions with energy gains of around 3–3.5 eV. Motivated by this finding, we have performed AIMD simulations with five different initial translational energies of CO molecules scattered off a 2O-MoSe2-vSe surface thermalized at 100 K. Our aim was to analyze how likely was the CO oxidation in this system and the corresponding reaction mechanisms. However, none of the 200 AIMD trajectories run for each energy has led to CO2 formation. The majority of the CO molecules are directly reflected, albeit in some cases (mainly at the lowest analyzed energies) some of them remain trapped at the surface without reacting with the adsorbed O atoms. The analysis of the trajectories show that those CO molecules impinging on the O are reflected or deviated to trapping before they can get close enough to the O atom in order recombination can occur. Inspection of the possible reaction paths suggests that the configurational space of the oxidation process via an Eley-Rideal process is very restrictive. Due to the unavoidable limited statistics and integration time, which does not allow to completely determine the final outcome of the trapped CO molecules, we can not absolutely exclude the possibility of the reaction to occur at a later stage. Nevertheless, we can safely conclude that, in spite of its large exothermicity, the recombinative CO+O abstraction is in the best case a very unlikely reaction in this system. As a general and important conclusion, the results of our dynamics simulations remark the importance of going beyond a pure static analysis of the reaction energetics to determine the catalytic character of any material.

### v2

@@ 4 Conclusions
We have studied the interaction of CO molecules coming from vacuum and a 2O-MoSe2-vSe surface, in which one O atom is adsorbed in a Se vacancy and a second O atom is adsorbed on top of a Se atom of the phase of the MoSe2 monolayer. The latter is the equilibrium structure upon (dissociative) adsorption of a O2 molecule on a MoSe2 monolayer with one Se vacancy 50. Interestingly, DFT calculations show that the CO recombination with the O atom adsorbed on top of the Se atom with either subsequent adsorption or desorption of the formed CO2 molecules are highly exothermic reactions with energy gains of around 3–3.5 eV. Motivated by this finding, we have performed AIMD simulations with five different initial translational energies of CO molecules scattered off a 2O-MoSe2-vSe surface thermalized at 100 K. Our aim was to analyze how likely was the CO oxidation in this system and the corresponding reaction mechanisms. However, none of the 200 AIMD trajectories run for each energy has led to CO2 formation. The majority of the CO molecules are directly reflected, albeit in some cases (mainly at the lowest analyzed energies) some of them remain trapped at the surface without reacting with the adsorbed O atoms. The analysis of the trajectories show that those CO molecules impinging on O are reflected or deviated to trapping before they can get close enough to the O atom in order recombination can occur. Inspection of the possible reaction paths suggests that the configurational space of the oxidation process via an Eley-Rideal process is very restrictive. Due to the unavoidable limited statistics and integration time, which does not allow to completely determine the final outcome of the trapped CO molecules, we can not absolutely exclude the possibility of the reaction to occur at a later stage. Nevertheless, we can safely conclude that, in spite of its large exothermicity, the recombinative CO+O abstraction is in the best case a very unlikely reaction in this system. As a general and important conclusion, the results of our dynamics simulations remark the importance of going beyond a pure static analysis of the reaction energetics to determine the catalytic character of any material.